Introduction

I saw an insect walking along the hallway in Building 76 this morning. “How fast is it moving?” I wondered. In this experiment, the speed of the insect will be determined using simple instruments.

1 Observations

I marked the hallway with pieces of tape, using a ruler to place each piece 5 cm from the previous one. Each time the insect’s forelegs reached a piece of tape, I noted the time with a hand-operated stopwatch.

Progress of the insect

Position (cm) Time (sec)
0  0.0
5  3.1
10  5.9
15  8.6
20  12.2

2 Analysis

A linear fit to the measurements yields

x = (0.598 ± 0.05 cm/s)t − 0.02 cm (1)

where x is the position of the insect and t the time.

3 Conclusion

This insect doesn’t walk as fast as a human.

Figure 1: Figure 1: Motion of an insect